Morning sessions

Monday 8:10-8:20

Opening ceremony 8:20-9:15 IPL

Hogben

9:15- 9:20 Mini-break

9:20-10:15 IPL

Jiang 10:15-10:45 Coffee break 10:45- 12:50

Parallel session 1 5 talks of 25min MS4, MS7, CT

12:50-15:30

Lunch break

Tuesday 8:10-9:05 IPL

Shaked-Monderer

9:05-9:10 Mini-break 9:10-10:05 IPL

Kressner

10:05-10:35 Coffee break 10:35-12:40

Parallel session 3 5 talks of 25min MS6, MS8, CT

12:40-15:30

Lunch break Wednesday 8:10-9:05 IPL

Peña

9:05-9:10 Mini-break 9:10-10:05 IPL

Nagy

10:05-10:35 Coffee break 10:35-12:40

Parallel session 5 5 talks of 25min

MS1, MS4, MS7, CT 12:40-13:30

Lunch break 13:30-19:00

Excursion to Tulum

Thursday 8:10-9:05 IPL

Fassbender

9:05-9:10 Mini-break 9:10-10:05 IPL

Van Dooren

10:05-10:35 Coffee break 10:35-12:40

Parallel session 6 5 talks of 25min MS1, MS5, CT 12:40-14:00

Lunch break

Friday 8:10-9:05 IPL

Rosenthal

9:05-9:10 Mini-break 9:10-10:05 IPL

Spitkovsy

10:05-10:35 Coffee break 10:35-12:40

Parallel session 8 5 talks of 25min MS2, MS3, CT 12:40-15:30

Lunch break

MS1  Combinatorial Matrix Theory  MS2  Eigenproblems: Theory and computation
MS3 Implem. and Appl. issues in reg. least squares MS4  Linear Algebra Education
MS5  Linear Algebra in Model Reduction  MS6  Matrix functions and matrix equations
MS7  Max Algebra  MS8 Nonnegative and eventually nonnegative matrices
CT  Contributed talks

Afternoon sessions

Monday

15:30-16:25 IPL

Boettcher

16:25-16:55 Coffee break 16:55-19:00

Parallel session 2 5 talks of 25min

MS8, CT

Tuesday

15:30-16:25 IPL

M. Dopico

16:25-16:55 Coffee break 16:55-19:00

Parallel session 4 5 talks of 25min MS4, MS6, MS7,CT Wednesday

Excursion

to Tulum Thursday 14:00-15:25 ILAS Business Meeting 15:30-16:25 IPL

Gemignani

16:25-16:55 Coffee break 16:55-19:00

Parallel session 7 5 talks of 25min MS2, MS3, CT

20:00 ILAS Banquet

Friday

15:30-17:35

Parallel session 9 5 talks of 25min MS1, MS5, CT

17:35-17:45

Closing ceremony

MS1  Combinatorial Matrix Theory  MS2  Eigenproblems: Theory and computation
MS3 Implem. and Appl. issues in reg. least squares MS4  Linear Algebra Education
MS5  Linear Algebra in Model Reduction  MS6  Matrix functions and matrix equations
MS7  Max Algebra  MS8 Nonnegative and eventually nonnegative matrices
CT  Contributed talks

Schedule of the Mini Symposia

MS1 Combinatorial Matrix Theory; Par. Sessions 5, 6, 9. MS2 Eigenproblems: Theory and computation; Par. Sessions 7, 8. MS3 Implementation and Application issues in regularizing least squares and total least squares; Par. Sessions 7, 8. MS4 Linear Algebra Education; Par. Sessions 1, 4, 5. MS5 Linear Algebra in Model Reduction; Par. Sessions 6, 9. MS6 Matrix functions and matrix equations; Par. Sessions 3, 4. MS7 Max Algebra; Par. Sessions 1, 4, 5. MS8 Nonnegative and eventually nonnegative matrices; Par. Sessions 2, 3.

Schedule of the Contributed talks

CT Alphabetic list of Contributed talks (pdf file).

Notes: IPL = Invited Plenary Lecture Plenary talks are 50 min plus 5 min for questions and comments, total=55 min. All other talks (MS and contributed) are 20 min. plus 5 min. for questions, total=25 min. In most Parallel sessions there will be 2 simultaneous CTs.

ILAS Student Luncheon, Tuesday, June 17, 12:40 hours. ILAS Business Meeting: Thursday, June 19, 14:00-15:25 hours. Banquet, Thursday, June 19, 20:00 hours.

Invited plenary lectures

• Albrecht Boettcher (Chemnitz, Germany)

• Froilán Dopico (Madrid, Spain)

• Heike Fassbender (Braunschweig, Germany)

• Luca Gemignani (Pisa, Italy)

• Leslie Hogben (Ames, Iowa, USA)

• Erxiong Jiang (Shanghai, China)

• Daniel Kressner (Zurich, Switzerland)

• James Nagy (the SIAM Lecture) (Atlanta, Georgia, USA)

• Juan Manuel Peña (the LAMA Lecture) (Zaragoza, Spain)

• Peter Rosenthal (Toronto, Canada)

• Naomi Shaked-Monderer (Emek Yezreel, Israel)

• Ilya Spitkovsky (Williamsburg, Virginia, USA)

• Paul Van Dooren (Louvain, Belgium)